OpenStudy (anonymous):

find cos(x) given than sin2x = -3/5, 3pi/2>2x>2pi

OpenStudy (anonymous):

no clue how to do this=p all help really appreciated

OpenStudy (anonymous):

so far im attempting to reason it out like this: sin2x=2sinxcosx soooo -3/5=2sinxcosx (-3/5)/2 = -3/10=sinxcosc

OpenStudy (anonymous):

You find what 2x is then divide your answers by 2 to find x.

OpenStudy (anonymous):

Find cos(2x) That'll get you on the right track :)

OpenStudy (anonymous):

\[\large \sin 2x=-\frac{3}{5}\] \[\large 2x=\sin^{-1} \frac{3}{5}\] The answers will be in the third and fourth quadrant.

OpenStudy (anonymous):

\[\huge \cos(2x) = \sqrt{1-\sin^2(2x)}\]

OpenStudy (anonymous):

And then divide by 2 to find x-calues.

OpenStudy (anonymous):

although I'm betting cos(2x) = 4/5 or -4/5

OpenStudy (anonymous):

but then again, you specified a quadrant, so cos(2x) is 4/5 after all? :D

OpenStudy (anonymous):

oh wait... it's -4/5

OpenStudy (anonymous):

man you guys are fast! thanksfor all the help so far!! working on finding cos2x

OpenStudy (anonymous):

Okay, I'm dizzy...

OpenStudy (anonymous):

haha! so i found that cos2x is 4/5, but are you allowed to divide cos2x by 2 in order to find cosx?

OpenStudy (anonymous):

\[\huge \frac{3\pi}{2}<2x<2\pi\] it must be this interval... then cos(2x) must be positive.

OpenStudy (anonymous):

or maybe its a formula i need to use=p gotta check

OpenStudy (anonymous):

\[\large 2x=216`52', 323`8'\] Using that restriction. \[\large 2x= 323`8'\] This ` symbol denotes a degree sign. Don't have it on my keyboard. Then \[\cos 2x=\cos 2(323`8')\]

OpenStudy (anonymous):

silly @mrdrew445 \[\huge \cos(x) =\sqrt{\frac{1+\cos(2x)}{2}}\] ^.^

OpenStudy (anonymous):

Half angle identities are the way to go ^.^

OpenStudy (anonymous):

ooooooo i didnt even realize! i didn't think i could put a double angle in there because it just says "1+cos(theta)" duhhhh lol

OpenStudy (anonymous):

Oops made an error \[\large \cos2x=\cos 323`8'\] And @PeterPan you're going along the long road. This question always gets students to solve it by using half angle identities etc. This is just using logic.

OpenStudy (anonymous):

and a calculator, @Azteck which is likely not at hand ^.^

OpenStudy (anonymous):

The answer is \[\frac{4}{5}\]

OpenStudy (anonymous):

That's only if it was cos2x. Your question says cosx

OpenStudy (anonymous):

^nope \[\huge \frac{3\pi}{2}<2x<2\pi\]\[\huge \frac{3\pi}{4}<x<\pi\] and cosine is negative in this interval ^.^

OpenStudy (anonymous):

You could get cos(x) without a calculator. I'm quite sure the teacher was shooting for the students to use identities on this one like I said\[\huge \cos(x) =\sqrt{\frac{1+\cos(2x)}{2}}\] and you already know what cos(2x) is

OpenStudy (anonymous):

sooo i ended up getting 3/5, but that doesn't seem right, gunna do it over to check

OpenStudy (anonymous):

thank you very much peter and azteck!!!

OpenStudy (anonymous):

just to make sure... cos(x) should be negative, ok? \[\huge \cos(x) =-\sqrt{\frac{1+\cos(2x)}{2}}\]

OpenStudy (anonymous):

because x ends up in an interval where cosine is negative

OpenStudy (anonymous):

Has he even learnt half angle identities. QUestions like this usually come apparaent once you learnt he first few identities.

OpenStudy (anonymous):

apparent*

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

i just recently learned half-angle identities, a few days ago

OpenStudy (anonymous):

There you have it. This is a logical test of knowledge... of half-angle identities ^.^

OpenStudy (anonymous):

Okay then Peter Pan's method is the way to go.

OpenStudy (anonymous):

To be honest, I've never learnt half angle results through my years. So my way was because I lacked knowledge of half angle results. Our syllabus does not require you to know half angle results.

OpenStudy (anonymous):

Half-angle identities are a pain ^.^ and not as useful as the others, it would seem :D

OpenStudy (anonymous):

all these identities and proving identities i need alot more work on! i find them all confusing=\ and im set to be taking calculus next year i must master them haha!

OpenStudy (anonymous):

@PeterPan That's basically why it isn't in our syllabus. LOL

OpenStudy (anonymous):

Yes~ you must^.^

OpenStudy (anonymous):

thanks peter appreciate it alot. fairly certain i have the right answer now (-3sqrt10/10)